# Variational Self-Supervised Learning

### v2

@@ III-G Implementation Details
$\epsilon\sim\left|\mathcal{N}(0,1)\right|,\quad z=\mu+\sigma\cdot\epsilon$
@@
Cosine-Based KL Divergence:
$\tilde{D}_{\mathrm{KL}}$ (${}^{\cos}(\mu_{1},\mu_{2},\sigma_{1}^{2},\sigma_{2}^{2})=\frac{1}{2}\Big{[}\log\big{(}1-\cos(\sigma_{1},\sigma_{2})\big{)}$
$+\big{(}\mathrm{softplus}_{\beta=3}\big{(}-\cos(\mu_{1},\mu_{2})\big{)}\big{)}^{2}$
$+\big{(}\mathrm{softplus}_{\beta=3}\big{(}-\cos(\sigma_{1},\sigma_{2})\big{)}-1\big{)}\Big{]}$
Expected Cosine-Based Log-Likelihood:
~$\tilde{\mathcal{L}}^{\cos}_{\mathrm{ll}}$ $(\mu_{1},\mu_{2},\sigma_{1}^{2},\sigma_{2}^{2})=\log\big{(}1-\cos(\sigma_{1},\sigma_{2})\big{)}$
$+4\,\mathrm{softplus}_{\beta=1}\big{(}-\cos(\sigma_{1},\sigma_{2})\big{)}$
$+\Big{[}\big{(}\mathrm{softplus}_{\beta=1}\big{(}-\cos(\mu_{1},\mu_{2})\big{)}\big{)}^{2}$
×$\qquad\times\mathrm{softplus}_{\beta=1}\big{(}-\cos(\sigma_{1},\sigma_{2})\big{)}\Big{]}$
The softplus function with temperature parameter $\beta$ is defined as:
$\mathrm{softplus}_{\beta}(x)=\frac{1}{\beta}\log\big{(}1+\exp(\beta x)\big{)}$
Cosine similarity provides a robust angular measure, particularly effective in high-dimensional embedding spaces, offering an alternative to Euclidean-based alignment in variational frameworks.
@@
IV Experimental Evaluation

### v3

@@ III-G Implementation Details
$\epsilon\sim\left|\mathcal{N}(0,1)\right|,\quad z=\mu+\sigma^{2}\cdot\epsilon$
@@
Let $S_{\beta}(a,b)=\mathrm{softplus}_{\beta}(-\cos(a,b))$, where $\cos(a,b)$ denotes the cosine similarity between vectors $a$ and $b$, and $\mathrm{softplus}_{\beta}$ is the softplus function scaled by the parameter $\beta$.
Cosine-Based KL Divergence: This formulation defines a divergence measure that replaces traditional Euclidean distance terms with cosine-based similarities, promoting angular alignment in latent distributions. The softplus-transformed cosine terms ensure smooth optimization, with a sharper focus on directional consistency between the mean and variance vectors.
$\tilde{D}_{\mathrm{KL}}^{\cos}(\mu_{1},\mu_{2},\sigma_{1}^{2},\sigma_{2}^{2})=\frac{1}{2}\Big{[}\log\big{(}S_{3}(\sigma_{1}^{2},\sigma_{2}^{2})\big{)}$
$+\big{(}S_{3}(\mu_{1},\mu_{2})\big{)}^{2}$
$+S_{3}(\sigma_{1}^{2},\sigma_{2}^{2})-1\Big{]}$ (9)
Expected Cosine-Based Log-Likelihood: The expected log-likelihood is reformulated to leverage cosine similarities, integrating both angular alignment of mean vectors and the spread (variance) of latent representations. This approach emphasizes semantic coherence between distributions, providing a robust alternative to standard log-likelihood terms in high-dimensional embedding spaces.
~$\tilde{\mathcal{L}}^{\cos}_{\mathrm{ll}}(\mu_{1},\mu_{2},\sigma_{1}^{2},\sigma_{2}^{2})=\log\big{(}S_{1}(\sigma_{1}^{2},\sigma_{2}^{2})\big{)}$
$+4\,S_{1}(\sigma_{1}^{2},\sigma_{2}^{2})$
$+\big{(}S_{1}(\mu_{1},\mu_{2})\big{)}^{2}\cdot S_{1}(\sigma_{1}^{2},\sigma_{2}^{2})$ (10)
Cosine similarity, as employed here, offers a resilient angular metric that remains effective in high-dimensional latent spaces. By substituting Euclidean-based alignment with cosine-driven measures, this framework enhances representational alignment and stability in variational inference models.
@@
Algorithm 1 outlines the training procedure for the proposed Variational Self-Supervised Learning (VSSL) framework. It iteratively updates the student and teacher networks by aligning their latent posterior distributions through KL divergence while optimizing an expected log-likelihood objective to enhance representation quality.